root(1-root(1-root(2-x2)))

Grade:

2 Answers

Pratham Ashish
14 years ago

hi ,

 for  root(2-x2) to be real we must have x^2=<2 emplies  that

x ε [-root2 , root 2]

then, for (1-root(2-x2) to be real , root(2-x2)=<1 , emplies (2-x2)=<1

x^2 >= 1  so , x ε [-root2 , -1 ] U [1, root2 ]

 we need not further analysis becoz   root(1-root(2-x2)) will always less than 1

so, the domain will be     x ε [-root2 , -1 ] U [1, root2 ]

 

RAHUL GAUR
13 years ago

-2<=x<=-1 and 1<=x<=2